What order of operations would I use to solve APR

     1.0646 = [1 + (APR)(92 / 365)]365 / 92 ;
Answer #1

1.0646 = [1 + (APR) * (92/365)] ^ (365/92)
1.0646 ^ (92/365) = 1 + (APR) * (92/365)
1.0646 ^ 0.252055 = 1 + (APR) * 0.252055
1.01590 = 1 + (APR) * 0.252055
1.01590 - 1 = (APR) * 0.252055
0.01590 = (APR) * 0.252055
0.01590 / 0.252055 = APR
APR = 0.0631 or 6.31%